OpenStudy (anonymous):

the letter PERIMETER are place in the bag ,you draw out one letter then you pick a second letter without replacing the first one ,how would you calculate the probability that both letters are vowles?

OpenStudy (anonymous):

Well, for the first one you have nine letters total and four vowels. For the second, you have eight letters total and three vowels left, assuming the first one was a vowel. So you just multiply those two probabilities together, \[ \frac{4}{9}\cdot\frac{3}8=\frac{1}{6} \]

OpenStudy (anonymous):

great ! ty